OpenStudy (tainted):

Solve and show work please :) -4√(x+9)=20

OpenStudy (raden):

-4sqrt(x+9)=20 or sqrt(x+9)=20/-4 sqrt(x+9)=-5 we knowed that, sqrt(x+9) would be real, if x+9>=0 or x>=-9 therefore, no solution for x satisfying to sqrt(x+9)=-5 !!